OpenStudy (anonymous):

plz any one tell me about domain and range of two variables functions ??????????

OpenStudy (anonymous):

plzzzzzzzz any one tell me

OpenStudy (amigatour):

The domain is x-variables and range is y-variables.

OpenStudy (anonymous):

yah yah

OpenStudy (amigatour):

The domain is an input and the range is an output.

OpenStudy (anonymous):

okz.......its clear but m confusig in two varialbe function in domain and range

OpenStudy (anonymous):

plzzzzzzzzzz reply me hurry up cox i'll submit my assignment tomarow

OpenStudy (amigatour):

(x,y) are the two variables. You shouldn't have procrastinated on this problem. It doesn't get easier.

OpenStudy (anonymous):

okz

OpenStudy (haseeb96):

consider a function \[f(x)=\sqrt{1+x^2+y^2}\] (x,y)=(2,3) here , the domain is (2,3) x=2 and y=3 now put it into the given equation \[f(x)=\sqrt{1+4+9}\] \[f(x)=\sqrt{14}\] here the \[\sqrt{14}\] is the range of the two variables

OpenStudy (amigatour):

I understand this a little. What I know is it's the f(x,y).

OpenStudy (anonymous):

Haseeb96 .... i under stand what u tell but if domain are not given then how we find domain and range of same question dt u r writeen in above

OpenStudy (amigatour):

You find the domain in x-values. Example (0,1). The domain is zero and the range is 1.

OpenStudy (anonymous):

its clear sis but if domain points r not given then how v find domain nd range

OpenStudy (amigatour):